OpenStudy (anonymous):

rewrite 2x+5y=10 in the slope intercept form

OpenStudy (gorv):

slope intercept theorm is y=mx+c so subtract 2x from both side

OpenStudy (gorv):

5y=-2x+10 now divide both side by 5

OpenStudy (anonymous):

2x - 5y = 10 -5y = -2x + 10 5y = 2x - 10 y = 2/5x - 2 Slope = 2/5 Y-intercept = -2